# Visualized Text-to-Image Retrieval

### v1

@@ Abstract
We propose Visualize-then-Retrieve (VisRet), a new paradigm for Text-to-Image (T2I) retrieval that mitigates the limitations of cross-modal similarity alignment of existing multi-modal embeddings. VisRet first projects textual queries into the image modality via T2I generation. Then, it performs retrieval within the image modality to bypass the weaknesses of cross-modal retrievers in recognizing subtle visual-spatial features. Experiments on three knowledge-intensive T2I retrieval benchmarks, including a newly introduced multi-entity benchmark, demonstrate that VisRet consistently improves T2I retrieval by 24.5% to 32.7% NDCG@10 across different embedding models. VisRet also significantly benefits downstream visual question answering accuracy when used in retrieval-augmented generation pipelines. The method is plug-and-play and compatible with off-the-shelf retrievers, making it an effective module for knowledge-intensive multi-modal systems. Our code and the new benchmark are publicly available at https://github.com/xiaowu0162/Visualize-then-Retrieve.
[…]
tcb@breakable
Visualized Text-to-Image Retrieval
Di Wuthanks:  Equal contribution., Yixin Wan11footnotemark: 1, Kai-Wei Chang University of California, Los Angeles {diwu,elaine1wan,kwchang}@cs.ucla.edu
@@ 1 Introduction
A common approach to T2I retrieval is to embed both the query and candidate images into a shared representation space, where similarity scores are computed (Frome et al., 2013; Kiros et al., 2014). However, obtaining accurate similarity rankings that capture fine-grained semantics in both text and image remains a long-standing challenge. Prior studies have observed that cross-modal embeddings often behave like “bags-of-concepts”, failing to model structured relationships among visual elements (Yüksekgönül et al., 2023; Kamath et al., 2023). For instance, Figure 1 presents a query that requires images of an entity (a Barnacle Goose) at specific postures (wings unfolded) to answer. While the embedding model succeeds at matching the entity type, it struggles to recognize subtler visual-spatial features such as the pose of the wing (unfolded) and the camera perspective (up-shot). To address these limitations, existing work has explored improving the embedding quality (Radford et al., 2021; Yu et al., 2022), query reformulation (Levy et al., 2023), and multi-stage reranking pipelines (Liu et al., 2024; Feng et al., 2025). Yet, all these strategies are ultimately constrained by the intrinsic difficulty of cross-modal similarity alignment, as they cannot bypass the stage of text-to-image similarity search.
@@
We propose Visualize-then-Retrieve (VisRet), a novel retrieval paradigm that decomposes T2I retrieval into two stages: text-to-image modality projection followed by within-modality retrieval. Concretely, the textual query is first visualized as one or more images via a T2I generation model. Then, the visualized query, which better exhibits the desired visual-spatial features, is used to perform image-to-image retrieval.
@@
Compared to prior methods, VisRet offers two key advantages. First, visualizations provide a more expressive and intuitive medium for encoding multiple compositional concepts such as entities, poses, and spatial relations, which are difficult to express via text alone. As shown in Figure 1, the visualized query is able to accurately depict the desired entity, posture, and camera angle at the same time. Second, by operating entirely within the image modality during retrieval, VisRet avoids the weaknesses of cross-modal retrievers and instead leverages the stronger capacity of these retrievers in uni-modal retrieval (Koishigarina et al., 2025).
@@
We evaluate VisRet on three challenging T2I retrieval benchmarks: INQUIRE-Rerank (Vendrow et al., 2024), Visual-RAG (Wu et al., 2025), and Visual-RAG-ME, a new benchmark we introduce that features feature comparison questions across multiple entities. Results show that VisRet substantially outperforms baseline T2I retrieval methods and a strong LLM-based query rewriting approach (section 4.2). When CLIP (Radford et al., 2021) is used as the retriever, VisRet outperforms the two baselines by 32.7% and 15.6% higher NDCG@10 respectively, averaged over three benchmarks. With E5-V (Jiang et al., 2024) as the retriever, the performance gain becomes 24.5% and 12.4%. Moreover, VisRet enhances downstream performance in retrieval-augmented generation (RAG) settings (section 4.3). It improves T2I question answering accuracy on Visual-RAG and Visual-RAG-ME by 3.8% and 15.7% in top-1 retrieval setting and 3.9% and 11.1% in top-10 retrieval setting.
@@ 2 Related Work
Early T2I retrieval benchmarks evaluate the ability to identify images based on their paired human-written captions. These datasets span multiple domains and include widely used benchmarks such as Flickr8K (Hodosh et al., 2013), Flickr30K (Young et al., 2014), and Fashion200K (Han et al., 2017). As multi-modal embedding models have matured, more challenging benchmarks have been introduced to assess retrieval in knowledge-intensive settings. These newer datasets—such as WebQA (Chang et al., 2022), INQUIRE (Vendrow et al., 2024), Visual-RAG (Wu et al., 2025), and MRAG-Bench (Hu et al., 2024)—shift the focus from caption matching to retrieving images that contain the knowledge necessary to answer complex natural language questions. These tasks challenge retrieval systems to support downstream reasoning in retrieval-augmented generation (RAG) pipelines.
@@
There have been extensive research on improving text-to-image retrieval from different perspectives. First, a number of works aim to train better multi-modal embeddings by designing better training objectives and data mixtures (Faghri et al., 2018; Radford et al., 2021; Yu et al., 2022; Li et al., 2022). Other studies improve various stages in the retrieval pipeline, such as textual query expansion (Levy et al., 2023; Lee et al., 2024) and reranking (Liu et al., 2024; Feng et al., 2025). Finally, a recent line of work introduces generative image retrieval (Li et al., 2024; Qu et al., 2025), which trains a generative model to directly memorize an index of the image corpus. Different from these existing approaches, VisRet expands the query semantics by directly visualizing it in the image space, thereby alleviating the workload of cross-modal retrieval. In addition, VisRet is a training-free plug-and-play framework that can accommodate any off-the-shelf retriever.
@@ 3 Approach
Retrieval Method Visual-RAG Visual-RAG-ME INQUIRE-Rerank-Hard
R@1 R@10 R@30 N@1 N@10 N@30 N@1 N@10 N@30  N@1   N@10 N@30
@@
Original Query 0.210 0.583 0.737 0.210 0.355 0.385 0.220 0.423 0.435 0.000 0.355 0.412
Query Expansion 0.238 0.586 0.737 0.238 0.360 0.395 0.410 0.575 0.572 0.136 0.349 0.407
Visualize-then-Retrieve 0.251 0.645 0.793 0.251 0.431 0.438 0.460 0.632 0.605 0.170 0.452 0.455
                    - multi-image 0.246 0.637 0.772 0.246 0.414 0.421 0.480 0.629 0.605 0.237 0.428 0.469
@@
Original Query 0.240 0.568 0.706 0.240 0.386 0.407 0.340 0.465 0.486 0.000 0.319 0.407
Query Expansion 0.223 0.560 0.719 0.223 0.368 0.391 0.460 0.569 0.566 0.170 0.367 0.412
Visualize-then-Retrieve 0.299 0.673 0.801 0.299 0.452 0.461 0.560 0.643 0.622 0.220 0.377 0.425
                   - multi-image 0.307 0.645 0.772 0.307 0.442 0.446 0.520 0.640 0.617 0.203 0.384 0.445
Table 1: Evaluation results across three T2I retrieval benchmarks using different retrieval strategies and retrievers. The best results in each column within each retriever group are boldfaced. R = Recall. N = NDCG.
3.1 Problem Formulation
Given a textual query $q$ and an image corpus $\mathcal{I}$, the task of Text-to-Image retrieval aims to retrieve $n\geq 1$ images $y_{1},...,y_{n}\in\mathcal{I}$ that best correspond to the semantics in $q$. Our paper further considers the task of Visual Question Answering (VQA), where the query is a knowledge-seeking question, with an expected answer $a$.
In this paper, we consider a basic retrieval-augmented generation (RAG) VQA pipeline: A multi-modal retriever $\mathcal{R}$ retrieves $k$ images from $\mathcal{I}$, denoted as $\{r_{1},...,r_{k}\}\equiv\mathcal{R}(q,\mathcal{I})\subseteq\mathcal{I}$. Then, a large vision-language model (LVLM) $\mathcal{M}$ directly generates the answer based on the question and the retrieval results $\mathcal{M}(q,\mathcal{R}(q,\mathcal{I}))$.
3.2 Visualize-then-Retrieve
We introduce Visualize-then-Retrieve (VisRet), a two-staged T2I retrieval pipeline that bridges the modality gap through modality projection. Figure 1 illustrates the pipeline with an intuitive example.
Modality Projection
The first stage of VisRet leverages a T2I generation system $\mathcal{T}$ to directly generate $m$ visualizations $\{v_{1},...,v_{m}\}\equiv\mathcal{T}(q)$. Empirically, we find it helpful to use an LLM within $\mathcal{T}$ to first rephrase $q$ into a T2I instruction $q^{\prime}$, before feeding into existing T2I generation models such as Stable Diffusion (Esser et al., 2024). To generate diverse $\{v_{1},...,v_{m}\}$, randomness can be injected either into $q^{\prime}$ or into the T2I generation.
Within-Modality Retrieval
In the second stage, VisRet performs retrieval within the image modality. Specifically, each synthesized image $v_{i}\in\{v_{1},...,v_{m}\}$ is independently used to retrieve a ranked list of images from the corpus:
$\mathcal{R}(v_{i},\mathcal{I})=[r_{1}^{(i)},...,r_{k}^{(i)}],$
To aggregate the $m$ separate retrieval results, we apply Reciprocal Rank Fusion (RRF) (Cormack et al., 2009). RRF assigns a fusion score to each candidate image $r$ based on its rank across $m$ lists:
where $\text{rank}_{i}(r)$ is the rank position of image $r$ in list $\mathcal{R}(v_{i},\mathcal{I})$, and $\lambda$ is a hyperparameter that controls the influence of lower-ranked items. The final top-$k$ retrieval result is formed by selecting the highest-scoring images according to $\text{score}_{\text{RRF}}(r)$.
Dataset Question Ground Truth Baseline Generated Image VisRet
@@ 3.2 Visualize-then-Retrieve
Table 2: Examples: VisRet improves retrieval by highlighting visual features implied by the textual query.
4 Results and Analyses

### v2

# VisRet: Visualization Improves Knowledge-Intensive Text-to-Image Retrieval

@@ Abstract
Text-to-image retrieval (T2I retrieval) remains challenging because cross-modal embeddings often behave as bags of concepts and underrepresent structured visual relationships such as pose and viewpoint. We propose Visualize-then-Retrieve (VisRet), a new paradigm for T2I retrieval that mitigates this limitation of cross-modal similarity alignment. VisRet first projects textual queries into the image modality via T2I generation. Then, it performs retrieval within the image modality to bypass the weaknesses of cross-modal retrievers in recognizing subtle visual-spatial features. Across four benchmarks (Visual-RAG, INQUIRE-Rerank, Microsoft COCO, and our new Visual-RAG-ME featuring multi-entity comparisons), VisRet substantially outperforms cross-modal similarity matching and baselines that recast T2I retrieval as text-to-text similarity matching, improving nDCG@30 by 0.125 on average with CLIP as the retriever and by 0.121 with E5-V. For downstream question answering, VisRet increases accuracy on Visual-RAG and Visual-RAG-ME by 3.8% and 15.7% in top-1 retrieval, and by 3.9% and 11.1% in top-10 retrieval. Ablation studies show compatibility with different T2I instruction LLMs, T2I generation models, and downstream LLMs. VisRet provides a practical and principled path that energizes further advances in vision-language retrieval. Our code and the new benchmark are publicly available at https://github.com/xiaowu0162/Visualize-then-Retrieve.
VisRet: Visualization Improves Knowledge-Intensive
Text-to-Image Retrieval
Yixin Wan1thanks:  Equal contribution., Di Wu [^1] footnotemark: 1, Kai-Wei Chang1 1University of California, Los Angeles {diwu,kwchang}@cs.ucla.edu
@@ 1 Introduction
A common approach to T2I retrieval is to embed both the text query and image candidates into a shared representation space, where similarity scores are computed (Frome et al., 2013; Kiros et al., 2014). However, obtaining accurate similarity rankings that capture fine-grained semantic alignments between text and image remains a long-standing challenge. Prior studies have observed that cross-modal embeddings often behave like “bags-of-concepts” for similarity matching, failing to model structured relationships among visual elements (Yüksekgönül et al., 2023; Kamath et al., 2023). For instance, Figure˜1 presents a query that requires images of an entity (a Barnacle Goose) at specific postures (wings unfolded) to answer. While the embedding model succeeds at matching the entity type, it struggles to recognize subtler visual-spatial features such as the pose of the wing (unfolded) and the camera perspective (upward-facing shot). To address these limitations, existing work has explored improving the embedding quality (Radford et al., 2021; Yu et al., 2022), query reformulation (Levy et al., 2023), and multi-stage reranking pipelines (Liu et al., 2024; Feng et al., 2025). Nevertheless, all these strategies are ultimately constrained by the intrinsic difficulty of cross-modal similarity alignment, as they cannot bypass the stage of text-to-image similarity search.
@@
We propose Visualize-then-Retrieve (VisRet), a novel retrieval paradigm that decomposes T2I retrieval into two stages: text-to-image modality projection followed by within-modality retrieval. Concretely, VisRet first utilizes a text-to-image generation (T2I generation) model to visualize the textual query, highlighting the key desired visual-spatial features. Then, the visualized query is used to perform direct image-to-image retrieval.
@@
Compared to prior methods, VisRet offers two key advantages. First, visualizations provide a more expressive and intuitive medium for representing multiple compositional concepts, including entities, poses, and spatial relations. These concepts are often difficult to convey through text alone, and attempting to encode them exhaustively in a text query can impair cross-modal matching performance due to limitations in embedding quality. In contrast, as shown in Figure˜1, the visualized query is able to accurately depict the desired entity, posture, and camera angle at the same time. Second, by operating entirely within the image modality during retrieval, VisRet avoids the weaknesses of cross-modal retrievers and instead leverages the stronger capacity of these retrievers in unimodal retrieval (Koishigarina et al., 2025).
@@
To investigate the effectiveness of VisRet, we conduct exhaustive experiments on two commonly used cross-modal embedding models using four benchmarks: Visual-RAG (Wu et al., 2025b), INQUIRE-Rerank (Vendrow et al., 2024), Microsoft COCO (Chen et al., 2015), and Visual-RAG-ME, a new benchmark we introduce that features challenging visual feature comparison questions across multiple entities. Results show that VisRet substantially outperforms baseline cross-modal similarity matching as well as three strong baseline methods that casts T2I retrieval to text-to-text similarity matching (section˜5.2). When CLIP (Radford et al., 2021) is used as the retriever, VisRet outperforms the baselines by 0.125 higher nDCG@30, averaged over four benchmarks. With E5-V (Jiang et al., 2024) as the retriever, the performance gain becomes 0.121. Moreover, VisRet enhances downstream performance in retrieval-augmented generation (RAG) settings (section˜5.3). It improves T2I question answering accuracy on Visual-RAG and Visual-RAG-ME by 3.8% and 15.7% in top-1 retrieval setting and 3.9% and 11.1% in top-10 retrieval setting. Further analyses show that VisRet is compatible with different T2I instruction generation and downstream question answering LLMs. In addition, the retrieval performance consistently improves with stronger T2I models. Finally, although they are effective in retrieving visual knowledge, the generated image queries themselves are generally insufficient as standalone knowledge to answer downstream questions.
@@ 2 Related Work
Early T2I retrieval benchmarks evaluate the ability to identify images based on their paired human-written captions. These datasets span multiple domains and include widely used benchmarks such as COCO (Chen et al., 2015), Flickr8K (Hodosh et al., 2013), Flickr30K (Young et al., 2014), and Fashion200K (Han et al., 2017). As multimodal embedding models have matured, more challenging benchmarks have been introduced to assess retrieval in knowledge-intensive settings. These newer datasets, such as WebQA (Chang et al., 2022), INQUIRE (Vendrow et al., 2024), Visual-RAG (Wu et al., 2025b), and MRAG-Bench (Hu et al., 2024), shift the focus from caption matching to retrieving images that contain the knowledge necessary to answer complex natural language questions. These tasks challenge T2I retrieval systems to support downstream reasoning in retrieval-augmented generation (RAG) pipelines. Further extending this stream of work, our Visual-RAG-ME benchmarks T2I retrieval for reasoning over the visual features across multiple entities.
@@
Extensive research has focused on improving text-to-image retrieval from different perspectives. First, a number of works aim to train better multimodal embeddings by designing better training objectives and data mixtures (Faghri et al., 2018; Radford et al., 2021; Yu et al., 2022; Li et al., 2022). Other studies improve various stages in the retrieval pipeline, such as textual query expansion (Levy et al., 2023; Lee et al., 2024) and reranking (Liu et al., 2024; Feng et al., 2025; Ding et al., 2025). Finally, a recent line of work introduces generative image retrieval (Li et al., 2024; Qu et al., 2025), which trains a generative model to directly memorize an index of the image corpus. Different from these existing approaches, VisRet expands the query semantics by directly visualizing it in the image space, thereby alleviating the workload of cross-modal retrieval. In addition, VisRet is a training-free plug-and-play framework that can accommodate any off-the-shelf retriever and pre-computed image embedding indices.
@@ 3 Approach
3.1 Problem Formulation
Given a textual query $q$ and an image corpus $\mathcal{I}$, the task of Text-to-Image retrieval aims to retrieve $n\geq 1$ images $y_{1},...,y_{n}\in\mathcal{I}$ that best correspond to the semantics of $q$. Our paper further considers the task of Visual Question Answering (VQA), where the query is a knowledge-seeking question, with an expected answer $a$.
In this paper, we consider a basic retrieval-augmented generation (RAG) VQA pipeline: A multimodal retriever $\mathcal{R}$ retrieves $k$ images from $\mathcal{I}$, denoted as $\{r_{1},...,r_{k}\}\equiv\mathcal{R}(q,\mathcal{I})\subseteq\mathcal{I}$. Then, a large vision-language model (LVLM) $\mathcal{M}$ directly generates the answer based on the question and the retrieval results $\mathcal{M}(q,\mathcal{R}(q,\mathcal{I}))$.
3.2 Visualize-then-Retrieve
We introduce Visualize-then-Retrieve (VisRet), a two-stage T2I retrieval pipeline that bridges the modality gap through modality projection. Figure˜1 illustrates the pipeline with an intuitive example.
Modality Projection
The first stage of VisRet leverages a T2I generation system $\mathcal{T}$ to directly generate $m$ visualizations $\{v_{1},...,v_{m}\}\equiv\mathcal{T}(q)$. Concretely, based on the original query, an LLM first operates in the text space to draft T2I instruction $q^{\prime}$, describing the images that likely satisfy the feature requirements implied in $q$. Next, we use an existing T2I generation method, such as Stable Diffusion (Esser et al., 2024), to project $q^{\prime}$ into a list of image queries $\{v_{1},...,v_{m}\},v_{i}\in\mathcal{I}$. To generate diverse images, randomness can be injected either into $q^{\prime}$ or into the T2I generation. In this work, we sample $m$ times with the same $q^{\prime}$.
Within-Modality Retrieval
In the second stage, VisRet performs retrieval exclusively within the image modality. First, each synthesized image $v_{i}\in\{v_{1},...,v_{m}\}$ is independently used to retrieve a ranked list of images from the corpus:
$\mathcal{R}(v_{i},\mathcal{I})=[r_{1}^{(i)},...,r_{k}^{(i)}],$
Then, we apply Reciprocal Rank Fusion (RRF) (Cormack et al., 2009) to aggregate the $m$ retrieval results. RRF assigns a fusion score to each candidate image $r$ based on its rank across $m$ lists:
where $\text{rank}_{i}(r)$ is the rank position of image $r$ in list $\mathcal{R}(v_{i},\mathcal{I})$, and $\lambda$ is a hyperparameter that controls the influence of lower-ranked items. The final top-$k$ retrieval result is obtained by selecting the highest-scoring images according to $\text{score}_{\text{RRF}}(r)$.
4 Visual-RAG-ME Benchmark
To facilitate the design of intelligent T2I retrieval systems, we introduce a multi-entity question answering and T2I retrieval dataset Visual-RAG-ME as an extension of Visual-RAG. Concretely, Visual-RAG-ME constructs questions that compare the visual features between the organism covered in Visual-RAG with another similar entity, for which we manually annotate a new set of retrieval labels. The Visual-RAG-ME annotation pipeline consists of three steps as described below.
Second Entity Identification
The goal of this step is to identify entities that are biologically close to the original entities in Visual-RAG so that plausible and challenging questions can be constructed. For this purpose, for each entity covered in Visual-RAG, we use BM25 (Robertson and Walker, 1994) to retrieve ten entities with the most similar scientific names.
Query Composition and Filtering
In this step, we (the authors) manually review all the 391 questions in Visual-RAG and attempt to construct a corresponding multi-entity question. A question is constructed when we can identify images for the second entity that clearly depict the same feature as in the positive images for the original entity in Visual-RAG. The questions we compose generally take a comparison style that asks whether the two organisms have the same feature or which organism features a more extreme stylistic feature (e.g., lighter coloration, smoother surface etc.). We were able to construct 82 multi-entity questions after this step. Then, a round of filtering is performed to remove (1) the questions that both GPT-4o and GPT-4o-mini can answer correctly without image information and (2) questions that cover overly similar topics. The final Visual-RAG-ME contains 50 high-quality multi-entity queries.
Retrieval Label Annotation and Balancing
For each question, we collect images of the second entity from the iNaturalist 2021 dataset (Horn et al., 2021) and annotate their retrieval labels. A positive label is assigned only if the image clearly displays the feature required to answer the question. For some of the questions, we find that a large number of positive images exist in the iNaturalist database. We therefore implement a filtering step where at most 50 positive images are kept for each entity. Table˜3 and Appendix Table˜8 show examples with their questions and ground-truth images.
In Table˜1, we present the basic statistics of Visual-RAG-ME. While it has slightly more positive per entity compared to Visual-RAG, our new benchmark is still challenging, with both CLIP and E5-V achieving a low Recall@1 due to its lengthy and knowledge-intensive queries.
VR VR-ME IR-Hard $C$-Hard
# Queries 391 50 59 500
||Query|| (word count) 18.5 25.1 6.0 12.2
# Images (per entity) 264 263 100 5000
# Positives (per entity) 14.3 20.8 12.5 1
CLIP Recall@1 0.210 0.220 0.000 0.000
E5-V Recall@1 0.240 0.340 0.000 0.000
Table 1: Dataset statistics and baseline performance. VR = Visual-RAG, IR = INQUIRE-Rerank, and $C$ = Microsoft COCO 2017.
Retrieval Method Visual-RAG Visual-RAG-ME INQUIRE-Rerank-Hard COCO-Hard
N@1 N@10 N@30 N@1 N@10 N@30 N@1 N@10 N@30 N@1 N@10 N@30
@@ 4 Visual-RAG-ME Benchmark
Original Query 0.210 0.355 0.385 0.220 0.423 0.435 0.000 0.355 0.412 0.000 0.017 0.042
LLM Rewriting 0.238 0.360 0.395 0.410 0.575 0.572 0.136 0.349 0.407 0.014 0.047 0.093
Corpus Captioning (BLIP) 0.105 0.217 0.271 0.210 0.354 0.371 0.136 0.319 0.401 0.054 0.122 0.153
VISA Reranking (top-30) 0.223 0.378 0.388 0.240 0.443 0.457 0.000 0.000 0.000 0.000 0.000 0.000
Captioning Reranking (top-30) 0.200 0.336 0.367 0.160 0.362 0.390 0.000 0.000 0.000 0.000 0.000 0.000
VisRet 0.251 0.431 0.438 0.460 0.632 0.605 0.170 0.452 0.455 0.034 0.072 0.108
VisRet - multi-image 0.246 0.414 0.421 0.480 0.629 0.605 0.237 0.428 0.469 0.038 0.069 0.106
@@
Original Query 0.240 0.386 0.407 0.340 0.465 0.486 0.000 0.319 0.407 0.000 0.163 0.178
LLM Rewriting 0.223 0.368 0.391 0.460 0.569 0.566 0.170 0.367 0.412 0.038 0.133 0.182
Corpus Captioning (BLIP) 0.156 0.294 0.319 0.240 0.405 0.436 0.119 0.342 0.398 0.064 0.158 0.196
VISA Reranking (top-30) 0.279 0.413 0.419 0.350 0.495 0.500 0.022 0.002 0.002 0.000 0.000 0.000
Captioning Reranking (top-30) 0.248 0.369 0.396 0.400 0.429 0.454 0.000 0.000 0.005 0.000 0.000 0.000
VisRet 0.299 0.452 0.461 0.560 0.643 0.622 0.220 0.377 0.425 0.042 0.173 0.205
VisRet - multi-image 0.307 0.442 0.446 0.520 0.640 0.617 0.203 0.384 0.445 0.048 0.168 0.206
Table 2: nDCG (N@k) across four T2I retrieval benchmarks using different retrieval strategies and retrievers. Within each retriever group, the best results per column are bolded. We use different colors to denote similarity matching in text-to-image, text-to-text, or image-to-image style. VisRet achieves the state-of-the-art across most evaluation settings, especially excelling in handling knowledge-intensive queries from Visual-RAG and Visual-RAG-ME.
Dataset Query Ground Truth Baseline Scores Visualized Query VisRet Scores
@@
COCO Plate with pizza, knife and fork laying on edge of plate with two glasses next to them. [Uncaptioned image] Rank:39, nDCG@10: 0.00 [Uncaptioned image] Rank:3, nDCG@10: 0.3937
Table 3: Examples: VisRet improves retrieval by highlighting visual features implied by the textual query. "Baseline" refers to cross-modal retrieval with the original query. CLIP embedding scores are used for this visualization.
5 Main Results